OpenStudy (blackops2luvr):

can someone give me the command for; 15 sqrt in green?

OpenStudy (anonymous):

\( \color{green}{\sqrt{something}}\)

OpenStudy (anonymous):

ops.. it's not showing XD

OpenStudy (blackops2luvr):

\(\Huge \color{orange}{\text{Practice~time!~<3}}\)

OpenStudy (e.mccormick):

\(\bbox[border:5px solid green,8pt,red]{\Large\color{green}{\sqrt{15}}}\)

OpenStudy (e.mccormick):

Lots of fun things with the MathJax \(\LaTeX\) support. They have a good deal of them turned on here: http://docs.mathjax.org/en/latest/tex.html \(\unicode[.55,0.05]{x22D6} \)

OpenStudy (unklerhaukus):

``` \[\color{green}{15\sqrt{something}}\] ```

hartnn (hartnn):

\(\color {green }{15\sqrt{in ~ green}}\) `\(\color {green }{15\sqrt{in ~ green}}\)`

hartnn (hartnn):

or id you mean this!? :P \(\color {green }{15~ sqrt ~ in ~ green}\) `\(\color {green }{15~ sqrt ~ in ~ green}\)`

OpenStudy (blackops2luvr):

\(\color {blue}{LaTeX~Is~FuN}\)